OpenStudy (anonymous):

show that cot theta+ tan theta = csc theta (sec theta) is an identity by transforming the left side to right side

OpenStudy (ipwnbunnies):

\[\cot(x) = \frac{ \cos(x) }{ \sin(x) }\]

OpenStudy (ipwnbunnies):

\[\tan(x) = \frac{ \sin(x) }{ \cos(x) }\]

OpenStudy (ipwnbunnies):

This one is also helpful: \[1 = \cos^{2}(x) + \sin^{2}(x)\]